OpenStudy (anonymous):

At present Alan is 10 years younger than Scott. Ten years from now, Scott' s age will be 3 times the present age of Alan. Find the present age of Alan and Scott. Please help! :D

OpenStudy (anonymous):

If answer is -5 please tell me

OpenStudy (anonymous):

When I worked it out Scott' s age was unborn (-5)

OpenStudy (anonymous):

Too bad! I have to go to sleep now so keep thinking I'll be back on tomorrow ok? Bye for now........

OpenStudy (anonymous):

it's wrong.......the difference is 10 years not 3 years.

OpenStudy (anonymous):

let's Alan's age be x then Scott's age=x+10 Scott's age after 10 years=x+10+10=x+20 3x=x+20 2x=20 x=10 Alan's age(present)=10 Scott's present age=20